OpenStudy (anonymous):

(4/5y^4)(-20y)(1/2y^5) step my step please thanks so much

OpenStudy (anonymous):

\[\frac{4}{5y^4}\times (-20y)\times\frac{1}{2y^5}\]

OpenStudy (curry):

but let me check

OpenStudy (curry):

i keep coming up with -80y

OpenStudy (curry):

but not 100% sure

OpenStudy (anonymous):

No good answer...

OpenStudy (anonymous):

I need it step by step please

OpenStudy (anonymous):

is the thing i wrote correct?

OpenStudy (anonymous):

i mean is that the problem? if so we can do it step by step

OpenStudy (anonymous):

Yes it is (:

OpenStudy (anonymous):

OK the best way to do this is by using the laws of exponents. For this problem I'm going to use two very important ones:\[a ^{n}*a ^{m} = a ^{n+m}\] & \[a ^{-n} = 1/a ^{n}\]. Therefore lets summarize the exercise as follows:\[(4/5)(1/y ^{4})(-20y)(1/2)(1/y ^{5})\] into\[(4/5)(-20)(1/2)(y)(y ^{-4})(y ^{-5})\] by the second law. Now using the first law we have \[(4/5)(-20)(1/2)(y ^{1-4-5})\] which is \[(4/5)(-20)(1/2)(y ^{-8})\] and simplifying the arithmetic we have \[(-80)(y ^{-8})\] which can also be expressed as \[-80/y ^{8}\] by the second law. Peace, Love and Happiness from Puerto Rico